OpenStudy (anonymous):

The quotient of a number and four is twelve less than the number. What is the number? A. –4 B. 9.6 C. 8 D. 16

OpenStudy (anonymous):

D. 4*4 is 16. 16-12 is 4.

OpenStudy (texaschic101):

x/4 = x - 12 (multiply equation by 4) x = 4x - 48 -3x = -48 x = 16 check... x/4 = x - 12 16/4 = 16 - 12 4 = 4 (correct) your answer is 16.....lahei is correct :)